A Bibliography of Publications of Karsten Strehl

Karsten Strehl

Tel: ?n/a? FAX: ?n/a?

E-mail: strehl@computer.org (Internet)

25 November 2011 Version 1.03

Abstract

This bibliography records publications of Karsten Strehl.

Title word cross-reference

Automata [Str98].

calculation [Str93]. CASE [Str96]. Checking [ST98b, ST98a, ST99]. Controlled [TTN^{+}98]. Cosimulation [Str96, Str97]. Coupling [Str96].

Design [Str96, Str97]. Diagram [ST98b, ST98a, Str98, ST99]. Diagrams [TTN^{+}98].

Environment [Str96, Str97].

Fields [Str93]. Flow [TTN^{+}98].

Hardware [STZE99a, STZE99b]. Hardware/Software [STZE99a, STZE99b].

Implementation [Str96, Str97]. Interval [ST98b, ST98a, Str98, ST99].

Kernel [Str97].

Machine [TTN^{+}98]. Model [ST98b, ST98a, ST99].

nets [ST99]. Networks [ST98a].

Optimistic [Str97].

particle [Str93]. Software [STZE99a, STZE99b]. Petri [ST99]. Predictive [Str97]. Process [ST98a]. program [Str93].

SCF [TTN^{+}98]. Scheduling [STZE99a, STZE99b]. SimBa [Str96, Str97]. simulation [Str93]. State [TTN^{+}98]. Statemate [Str96]. Symbolic [ST98b, ST98a, Str98, ST99, STZE99a, STZE99b]. Synchronization [Str97]. Systems [STZE99a, STZE99b].

Techniques [ST98b, ST98a, Str98, ST99, STZE99a, STZE99b]. Timed [Str98]. Tool [Str96]. trajectories [Str93].

Using [ST98b, ST98a, Str98, STZE99a, STZE99b].

Verification [Str98].

References

Strehl:1998:SMCb

[ST98a] Karsten Strehl and Lothar Thiele.

Symbolic model checking of process networks using interval diagram techniques. In Proceedings of the IEEE/ACM International Conference on Computer-Aided Design (ICCAD-98), pages 686–692. San Jose, California, November 8–12, 1998.

Strehl:1998:SMCa

[ST98b] Karsten Strehl and Lothar Thiele.

Symbolic model checking using interval diagram techniques. Technical Report 40, Computer Engineering and Networks Lab (TIK), Swiss Federal Institute of Technology (ETH) Zurich, Gloriastrasse 35, CH-8092 Zurich, February 1998.

Strehl:1999:IDT

[ST99] Karsten Strehl and Lothar Thiele.

Interval diagram techniques for symbolic model checking of Petri nets. In Proceedings of the Design, Automation and Test in Europe Conference (DATE99). Munich, Germany, March 9–12, 1999.

Strehl:1993:FEP

[Str93] Karsten Strehl. Fields. A program for the calculation and simulation of particle trajectories. CoMet, Publishers for Educational Software, Duisburg, later: Cornelsen Software, Berlin, 1993. ISBN 3-89418-717-4.

Strehl:1996:DIC

[Str96]  Karsten Strehl.  Design and

implementation of the coupling of the CASE tool Statemate to the cosimulation environment SimBa. Term paper, Electronic Systems and Microsystems (ESM) Research Department, Computer Science Research Center (FZI), Karlsruhe, and Institute for Information Processing Technology (ITIV), University of Karlsruhe, July 1996.

Strehl:1997:DIP

[Str97]  Karsten Strehl.  Design and

implementation of a predictive, optimistic synchronization kernel for the cosimulation environment SimBa. Diploma thesis, Electronic Systems and Microsystems (ESM) Research Department, Computer Science Research Center (FZI), Karlsruhe, and Institute for Information Processing Technology (ITIV), University of Karlsruhe, March 1997.

Strehl:1998:UID

[Str98] Karsten Strehl. Using interval diagram techniques for the symbolic verification of timed automata. Technical Report 53, Computer Engineering and Networks Lab (TIK), Swiss Federal Institute of Technology (ETH) Zurich, Gloriastrasse 35, CH-8092 Zurich, July 1998.

Strehl:1999:SHSa

[STZE99a] Karsten Strehl, Lothar Thiele, Dirk Ziegenbein, and Rolf Ernst. Scheduling hardware/software systems using symbolic techniques. Technical Report TIK-67, Computer Engineering and Networks Lab (TIK), Swiss Federal Institute of Technology (ETH) Zurich, Gloriastrasse 35, CH-8092 Zurich, January 1999.

Strehl:1999:SHSb

[STZE99b] Karsten Strehl, Lothar Thiele, Dirk Ziegenbein, and Rolf Ernst. Scheduling hardware/software systems using symbolic techniques. In Proceedings of the 7th International Workshop on Hardware/Software Codesign (CODES’99), Rome, Italy, 3–5 May, 1999, page ?? ????, ????, 1999. ISBN ???? LCCN ????

Thiele:1998:SSM

[TTN^{+}98] Lothar Thiele, Jürgen Teich, Martin Naedele, Karsten Strehl, and Dirk Ziegenbein. SCF—state machine controlled flow diagrams. Technical Report 33, Computer Engineering and Networks Lab (TIK), Swiss Federal Institute of Technology (ETH) Zurich, Gloriastrasse 35, CH-8092 Zurich, January 1998.